OpenStudy (anonymous):

A crane lifts a mass of 500 kg uniformly accelerated with a = 0.17 m/s^2 . What is the force on the crane's hook ?

OpenStudy (anonymous):

5750N m(g+a)=500(9.8+1.7)=5750N

OpenStudy (anonymous):

Negative or positive ?

OpenStudy (anonymous):

@zaphodplaysitsafe

OpenStudy (anonymous):

Hi @ali_rami1990, the above answer is correct. The hook exerts a force of 5750N on the mass to lift it up. Similarly the mass(or the connecting rope) pulls the hook downwards with the same force 5750N. here the question of negative or positive will not arise, as in the context of the question, it is just being asked about the force on the hook. Now the force on the hook is downwards. Here, we are not given any convention like up is +ve and down is -ve. so we can't define the sign here. but we know that the force of 5750N is pulling the hook downwards.

OpenStudy (anonymous):

The above answer is incorrect. The acceleration is 0.17 m/s^2, not 1.7 m/s^2... It should be 4985 N. However, like mentioned, the directional convention was not specified.

OpenStudy (anonymous):

@Pompeii00, thanks for cross-checking, I didn't really check the numerical calcualtion done above.

OpenStudy (anonymous):

sorry for my mistake in numericals but i guess my concept is correct

OpenStudy (anonymous):

I got it, thanks @Dhawal @zaphodplaysitsafe @Pompeii00 The best response for me is zaphodplaysitsafe's because he explained the concept, however Dhawal solved it as well, and Pompeii00 checked the answer.